# Granger Causality in High-Dimensional Networks of Time Series

### v3

@@ 1 Introduction
In fact, utilizing conventional bivariate vector autoregressive models (VAR) can be a practical and manageable way to identify pairwise causal interactions in high-dimensional networks as long as the potential confounding effect due to the high-dimensional background nodes is addressed. Adopting this strategy, we propose a procedure that uses conventional Granger causality (GC) testing to detect causal connections between any two nodes while accounting for the global background activity in a low-dimensional representation. Such a representation is constructed through spectral dynamic principal component analysis (sDPCA) (Brillinger, 1969, 1981), which extends the principal component analysis to incorporate temporal dependence. By working with the spectral density matrix —directly linked to all auto and cross-covariances- sDPCA preserves frequency-specific features critical for neural data analysis.
@@
The sDPCA enhances the PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining temporal information. In other words, while the conventional PCA is based on the eigendecomposition at the zero-lag (or contemporaneous) covariance matrix, sDPCA is based on the eigendecomposition of the spectral density matrix estimated for lagged auto-covariances at all frequencies. Since there is a 1-1 relationship between the spectral matrix (at all frequencies) and the covariance matrix (at all lags), this approach captures lagged cross-correlations amongst signals.
@@
Various methodologies have been derived from frequency domain PCA. For instance, Stoffer et al. (1993); Stoffer (1999); Stoffer et al. (2000) developed a spectral envelope using dynamic PCA to identify common signals in multiple categorical and real-valued time series, Hörmann et al. (2015) proposed a variant of dynamic PCA for analyzing functional time-series data. As demonstrated by Ombao and Ho (2006); Wang et al. (2019); Ombao and Pinto (2022), principal component analysis in the frequency domain offers practical benefits and is highly effective in summarizing high-dimensional signals by capturing their main modes of variation. On the other hand, given the multivariate nature of the objectives outlined herein, it becomes relevant to inquire about the applicability of the state-of-the-art techniques of multivariate time series analysis. For example, Factor-augmented VAR (FAVAR) models (Bernanke et al., 2005; Lin and Michailidis, 2020) reduce dimensionality by extracting a handful of static factors from large panels and then placing these factors into a more manageable VAR system. This strategy centers on global modeling and treating the extracted factors as regressors to explain multiple target variables; however, it is not designated for cases where attention is on a single pair of nodes within a high-dimensional network. Dynamic factor models (DFMs) (Forni et al., 2000, 2015; Stock and Watson, 2005; Barigozzi and Hallin, 2017) similarly rely on a few latent components, focusing on decomposing each series into common and idiosyncratic components. DFM allows more detailed lagged structures than static factor approaches but often seeks to explain broad data variance or perform large-scale forecasts. Conditional Granger causality (CGC) (Geweke, 1984) instead tests whether adding one predictor refines forecasts of another, holding all others fixed. Although CGC reveals causal associations, it can become complicated in high dimensions when many variables must be jointly conditioned. While there are some parallels between the proposed approach and that of others, each method is characterized by a distinct emphasis. The proposed approach diverges by isolating specific node pairs. It accounts for background effects by spectral components and then applies GC testing on a simplified bivariate system. This diverges from CGC’s exhaustive conditioning on many channels, from FAVAR’s global factor plus VAR, and dynamic factor models’ emphasis on large-scale shared patterns. The approach also provides feasibility in large networks, where a smaller factor-driven representation is sufficient to identify causal interactions.
@@
The proposed approach distinguishes itself from existing approaches by focusing on frequency-domain dimensionality reduction combined with targeted pairwise causality analysis. The theoretical formulations of conditional GC, FAVAR, and DFMs have different aims – identifying influences through conditioning, incorporating large information sets into VARs, and explaining common variance in high dimensions, respectively – and each comes with certain assumptions about the data-generating process. The proposed approach does not compete with these state-of-the-art methods on their own terms; rather, it occupies a complementary cavity. It assumes that in many high-dimensional systems (e.g., brain recordings), one can extract the key oscillatory components and concentrate on their interactions, thereby enabling meaningful causal understanding without having to model every other interaction. Empirically, this translates to a convenient method that is sensitive to the features of brain monitoring data – handling many channels and frequency-specific processes – while building on the proven strengths of factor modeling and Granger causality. Given the objectives of established methods and our approach, it is important to note that utilizing GC with sDPCA does not propose a universal substitute for full network modeling. Instead, it is a domain-orientated alternative that emphasizes interpretability and feasibility in the analysis of high-dimensional network time series. This allows for targeted analysis, whether at the level of individual node pairs or among a tractable number of node clusters, without being constrained by finitely many interactions. In what follows, Section 2 presents the proposed approach for an HDN setting, Section 3 describes simulation setups and results, and Section 4 features an application to EEG recordings, showing how the method handles brain signals.

### v4

@@ 1 Introduction
In fact, utilizing conventional bivariate vector autoregressive models (VAR) can be a practical and manageable way to identify pairwise causal interactions in high-dimensional networks as long as the potential confounding effect due to the high-dimensional background nodes is addressed. Adopting this strategy, we propose a procedure that uses conventional Granger causality (GC) testing to detect causal connections between any two nodes while accounting for the global background activity in a low-dimensional representation. Such a representation is constructed through spectral dynamic principal component analysis (sDPCA) (Brillinger, 1969, 1981), which extends the principal component analysis to incorporate temporal dependence. By working with the spectral density matrix, directly linked to all auto and cross-covariances, sDPCA preserves frequency-specific features critical for neural data analysis.
@@
The sDPCA enhances the PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining temporal information. In other words, while the conventional PCA is based on the eigendecomposition at the zero-lag (or contemporaneous) covariance matrix, sDPCA is based on the eigendecomposition of the spectral density matrix estimated for lagged auto-covariances at all frequencies. Since there is a 1-1 relationship between the spectral matrix (at all frequencies) and the covariance matrix (at all lags), this approach captures lagged cross-correlations amongst signals. From a Bayesian paradigm, the generalized probabilistic PCA developed by Gu and Shen (2020) provides a dimension-reduction mechanism that accounts for serial and spatial correlation in multivariate signals.
@@
Various methodologies have been derived from frequency domain PCA. For instance, Stoffer et al. (1993); Stoffer (1999); Stoffer et al. (2000) developed a spectral envelope using dynamic PCA to identify common signals in multiple categorical and real-valued time series, Hörmann et al. (2015) proposed a variant of dynamic PCA for analyzing functional time-series data. As demonstrated by Ombao and Ho (2006); Wang et al. (2019); Ombao and Pinto (2022), principal component analysis in the frequency domain offers practical benefits and is highly effective in summarizing high-dimensional signals by capturing their main modes of variation. On the other hand, given the multivariate nature of the objectives outlined herein, it becomes relevant to inquire about the applicability of the state-of-the-art techniques of multivariate time series analysis. For example, Factor-augmented VAR (FAVAR) models (Bernanke et al., 2005; Lin and Michailidis, 2020) reduce dimensionality by extracting a handful of static factors from large panels and then placing these factors into a more manageable VAR system. This strategy centers on global modeling and treating the extracted factors as regressors to explain multiple target variables; however, it is not designated for cases where attention is on a single pair of nodes within a high-dimensional network. Dynamic factor models (DFMs) (Forni et al., 2000, 2015; Stock and Watson, 2005; Barigozzi and Hallin, 2017) similarly rely on a few latent components, focusing on decomposing each series into common and idiosyncratic components. DFM allows more detailed lagged structures than static factor approaches, but often seeks to explain broad data variance or perform large-scale forecasts. Bayesian mixture network models can likewise account for cross-graph heterogeneity (Yin et al., 2022). Conditional Granger causality (CGC) (Geweke, 1984) instead tests whether adding one predictor refines forecasts of another, holding all others fixed. Although CGC reveals causal associations, it can become complicated in high dimensions when many variables must be jointly conditioned. While there are some parallels between the proposed approach and that of others, each method is characterized by a distinct emphasis. The proposed approach diverges by isolating specific node pairs. It accounts for background effects by spectral components and then applies GC testing on a simplified bivariate system. This diverges from CGC’s exhaustive conditioning on many channels, from FAVAR’s global factor plus VAR, and dynamic factor models’ emphasis on large-scale shared patterns. The approach also provides feasibility in large networks, where a smaller factor-driven representation is sufficient to identify causal interactions.
@@
The proposed approach distinguishes itself from existing approaches by focusing on frequency-domain dimensionality reduction combined with targeted pairwise causality analysis. The theoretical formulations of CGC, FAVAR, and DFMs have different aims – identifying influences through conditioning, incorporating large information sets into VARs, and explaining common variance in high dimensions, respectively – and each comes with certain assumptions about the data-generating process. The proposed approach does not compete with these state-of-the-art methods on their own terms; rather, it occupies a complementary cavity. It assumes that in many high-dimensional systems (e.g., brain recordings), one can extract the key oscillatory components and concentrate on their interactions, thereby enabling meaningful causal understanding without having to model every other interaction. Empirically, this translates to a convenient method that is sensitive to the features of brain monitoring data – handling many channels and frequency-specific processes – while building on the proven strengths of factor modeling and Granger causality. Given the objectives of our approach, it is important to note that GC combined with sDPCA does not propose a universal substitute for full network modeling. Instead, it is a domain-orientated alternative that emphasizes interpretability and feasibility in the analysis of high-dimensional network time series. This allows for targeted analysis, whether at the level of individual node pairs or among a tractable number of node clusters, without being constrained by finitely many interactions. In what follows, Section 2 presents the proposed approach for an HDN setting, Section 3 describes simulation setups and results, and Section 4 features an application to EEG recordings, showing how the method handles brain signals.